# Facilitating Battery Swapping Services for Freight Trucks with Spatial-Temporal Demand Prediction

### v1

@@ 1 Introduction
The global challenge of carbon emissions has exhibited different patterns and witnessed various responses across sectors. In 2021, 21.2% of these emissions are attributed to the transportation sector (Bouckaert et al. (2021)). While substantial progress has been made in the electrification of passenger vehicles, heavy-duty trucks remain notably falling behind in this transition. In contrast, heavy-duty trucks, though accounting for less than 10% of the total vehicle population, are responsible for more than 40% of carbon dioxide emissions (Hao et al. (2019); Moultak et al. (2017); Xue et al. (2019)). This disparity underscores the opportunity for mitigating transportation environment impact via the electrification of heavy-duty trucks. In practice, several factors contribute to the lagging electrification of heavy-duty trucks, chiefly among them are the limitations in battery size and energy density, range anxiety, and long charging durations. However, as EV technology advances and battery costs decrease, battery swapping stations are emerging as an economically viable solution for business proprietors and station operators (Zhu et al. (2023)).
@@
Our prediction approach is closely related to the literature on machine learning methods for traffic forecasting. As neural network (NN) methods evolve (Hochreiter and Schmidhuber (1997); Krizhevsky et al. (2012); Kipf and Welling (2016); Vaswani et al. (2017)), NN-based traffic prediction methods (Park and Rilett (1999); Fu et al. (2016); Zhao et al. (2019); Wang et al. (2020)) have become increasing effective. In the traffic prediction literature, there is also a growing interest in leveraging machine methods for climate-change-related goals (Blattner et al. (2021); Rollend et al. (2022); Buechler et al. (2021); Kapoor (2020)). The predicted traffic volume is used to optimize the operations of BSS, which is related to the vast literature of planning and management for electric vehicles (Mak et al. (2013); Widrick et al. (2018); Koirala et al. (2022); Qi et al. (2023); Sarker et al. (2014)) and energy systems (Yang et al. (2022); Powell et al. (2022); Borlaug et al. (2021)).
Here we summarize our key findings. From a prediction perspective, based on our dataset, we observe that the integration of the attention module (Vaswani et al. (2017)) enhances the prediction accuracy of the GCN-based model for longer horizons. In terms of the service rate, compared to the hindsight optimal strategy where the ground truth is known, our predict-then-optimize approach demonstrates robust performance across various environments. Our investigation also sheds light on the planning choice between fixed BSS (a BSS with a fixed location) and mobile BSS (a BSS that can change its location). We observe that the traffic pattern significantly influences decision-making for infrastructure choices more than the accuracy of traffic prediction: in the initial phases of implementing battery-swapping services, the traffic has a more irregular pattern/heteroskedasticity, which favors the mobile BSS; as the system matures, the traffic has lower variation and more stable pattern, and in this case, the fixed BSS is more preferred.
@@ 2.2 Model
  • Temporal Graph Convolutional Network (T-GCN) ((Zhao et al., 2019)): T-GCN combines the graph convolutional network (GCN) and the gated recurrent unit (GRU) to incorporate the spatial and temporal information to forecast traffic volumes.
  • Attention Graph Convolutional Network (A3T-GCN)((Bai et al., 2021)): A3T-GCN modifies the T-GCN by incorporating an attention module to dynamically capture the spatial and temporal correlations of traffic volumes.
@@
Now we describe a scheduling optimization model that dispatches these mobile batteries among service areas. In our highway network, the travel time for heavy-duty trucks between adjacent battery swapping stations (approximately 40 miles) is approximately 1 hour. As of the year 2023, one 282 kWh battery for heavy-duty trucks also takes roughly 1 hour to get fully charged with an ultra-fast charger ((Saadaoui et al., 2023)). Henceforth, we discretize time into hourly intervals, with a unit time step representing 1 hour. The optimization model needs to decide whether each mobile BSS should remain at the current station or be dispatched and relocated to a nearby station. If the latter option is adopted, the model needs to further decide the number of batteries to dispatch. The moving battery on the way cannot be used for service until it arrives at the destination.
@@ Appendix A Details for the Optimization Model
We schedule these mobile batteries in a rolling manner: in each time step, we solve the optimization problem to obtain a scheduling plan for the next $T$ hours, but only the scheduling plan for the current (first) time step is executed. We assume that there is a sufficient number of trucks serving as mobile BSS, ensuring that at every time step, we can deliver any battery in the mobile BSS to any nearby stations. We outline our model in model (A) associated with notations defined in Table 1.
@@
$\mathcal{S}$
the set of service stations/areas
$\mathcal{N}(i)$
the neighboring stations of station $i$
@@
$T$
the planning horizon
$Q_{i}$
the initial number of mobile batteries at station $i$
$F_{i}$
the number of fixed batteries at station $i$
$D_{i,t}$
battery demand (prediction) at station $t$ at time step $t$
@@
$z_{t,i,j}$
the number of batteries moving from station/area $i$ to $j$ at time step $t$
$L_{t,i}$
the number of lost/unmet demand at station $i$ at time step $t$
Constraints (1b) and (1c) are flow conservation constraints for the mobile batteries among stations. Constraint (1d) only allows mobile BSS to transfer to adjacent stations in consecutive time steps. Constraints (1e), (1f), and (1h) describe the definition of lost demand. The total number of batteries equals inventory level $\times$ average demand of all stations. The ratio of the total number of mobile/movable batteries and fixed/immovable batteries is set as the given mobile-to-fix ratio. At each station, the number of fixed batteries $F_{i}$ is set proportional to the average demand of station $i$. In the experiment, the initial value of $Q_{i}$ is determined by solving an optimization model similar to (A) with mobile battery total capacity constraint, which is formally written as
@@ Appendix B Other Graphs
Figure 3: For 2(a), the horizontal axis stands for the ratio of the total inventory of batteries to the average demand, and the vertical axis stands for the relative lost demand compared to that of the oracle (hindsight) scheduling policy. For 2(b), the horizontal axis stands for the value of $h$. Based on different $h$, the scheduling policy outputs different unmet demand, and the vertical axis depicts the ratio of unmet demand to the total demand.

### v2

@@ 1 Introduction
The global challenge of carbon emissions has exhibited different patterns and witnessed various responses across sectors. In 2021, 21.2% of these emissions are attributed to the transportation sector ([4]). While substantial progress has been made in the electrification of passenger vehicles, heavy-duty trucks remain notably falling behind in this transition. In contrast, heavy-duty trucks, though accounting for less than 10% of the total vehicle population, are responsible for more than 40% of carbon dioxide emissions ([7, 14, 24]). This disparity underscores the opportunity for mitigating transportation environment impact via the electrification of heavy-duty trucks. In practice, several factors contribute to the lagging electrification of heavy-duty trucks, chiefly among them are the limitations in battery size and energy density, range anxiety, and long charging durations. However, as EV technology advances and battery costs decrease, battery swapping stations are emerging as an economically viable solution for business proprietors and station operators ([27]).
@@
Our prediction approach is closely related to the literature on machine learning methods for traffic forecasting. As neural network (NN) methods evolve ([8, 12, 10, 21]), NN-based traffic prediction methods ([15, 6, 26, 22]) have become increasing effective. In the traffic prediction literature, there is also a growing interest in leveraging machine methods for climate-change-related goals ([2, 18, 5, 9]). The predicted traffic volume is used to optimize the operations of BSS, which is related to the vast literature of planning and management for electric vehicles ([13, 23, 11, 17, 20]) and energy systems ([25, 16, 3]).
Here we summarize our key findings. From a prediction perspective, based on our dataset, we observe that the integration of the attention module ([21]) enhances the prediction accuracy of the GCN-based model for longer horizons. In terms of the service rate, compared to the hindsight optimal strategy where the ground truth is known, our predict-then-optimize approach demonstrates robust performance across various environments. Our investigation also sheds light on the planning choice between fixed BSS (a BSS with a fixed location) and mobile BSS (a BSS that can change its location). We observe that the traffic pattern significantly influences decision-making for infrastructure choices more than the accuracy of traffic prediction: in the initial phases of implementing battery-swapping services, the traffic has a more irregular pattern/heteroskedasticity, which favors the mobile BSS; as the system matures, the traffic has lower variation and more stable pattern, and in this case, the fixed BSS is more preferred.
@@ 2.2 Model
  • Temporal Graph Convolutional Network (T-GCN) ([26]): T-GCN combines the graph convolutional network (GCN) and the gated recurrent unit (GRU) to incorporate the spatial and temporal information to forecast traffic volumes.
  • Attention Graph Convolutional Network (A3T-GCN)([1]): A3T-GCN modifies the T-GCN by incorporating an attention module to dynamically capture the spatial and temporal correlations of traffic volumes.
@@
Now we describe a scheduling optimization model that dispatches these mobile batteries among service areas. In our highway network, the travel time for heavy-duty trucks between adjacent battery swapping stations (approximately 40 miles) is approximately 1 hour. As of the year 2023, one 282 kWh battery for heavy-duty trucks also takes roughly 1 hour to get fully charged with an ultra-fast charger ([19]). Henceforth, we discretize time into hourly intervals, with a unit time step representing 1 hour. The optimization model needs to decide whether each mobile BSS should remain at the current station or be dispatched and relocated to a nearby station. If the latter option is adopted, the model needs to further decide the number of batteries to dispatch. The moving battery on the way cannot be used for service until it arrives at the destination.
@@ 3 Results
Acknowledgments and Disclosure of Funding
This work utilizes traffic data provided by Chongqing Expressway Group Company. The authors would like to thank Mengxiong Zhou, Jiahong Li, and Jie Yang for their collaboration on this project.
This work was partially supported by the National Natural Science Foundation of China under Grant 61936009, and partially by the National Science and Technology Innovation 2030 Major Project of the Ministry of Science and Technology of China (NO.2018AAA0101604).
@@ Appendix A Details for the Optimization Model
We schedule these mobile batteries in a rolling manner: in each time step, we solve the optimization problem to obtain a scheduling plan for the next $T$ hours, but only the scheduling plan for the current (first) time step is executed. We assume that there is a sufficient number of trucks serving as mobile BSS, ensuring that at every time step, we can deliver any battery in the mobile BSS to any nearby stations. We outline our model in model (1) associated with notations defined in Table 1.
@@
$\mathcal{S}$ the set of service stations/areas
$\mathcal{N}(i)$ the neighboring stations of station $i$
@@
$T$ the planning horizon
$Q_{i}$ the initial number of mobile batteries at station $i$
$F_{i}$ the number of fixed batteries at station $i$
$D_{i,t}$ battery demand (prediction) at station $t$ at time step $t$
@@
$z_{t,i,j}$ the number of batteries moving from station/area $i$ to $j$ at time step $t$
$L_{t,i}$ the number of lost/unmet demand at station $i$ at time step $t$
Constraints (1b) and (1c) are flow conservation constraints for the mobile batteries among stations. Constraint (1d) only allows mobile BSS to transfer to adjacent stations in consecutive time steps. Constraints (1e), (1f), and (1h) describe the definition of lost demand. The total number of batteries equals inventory level $\times$ average demand of all stations. The ratio of the total number of mobile/movable batteries and fixed/immovable batteries is set as the given mobile-to-fix ratio. At each station, the number of fixed batteries $F_{i}$ is set proportional to the average demand of station $i$. In the experiment, the initial value of $Q_{i}$ is determined by solving an optimization model similar to (1) with mobile battery total capacity constraint, which is formally written as
@@ Appendix B Other Graphs
Figure 4: For 4(a), the horizontal axis stands for the ratio of the total inventory of batteries to the average demand, and the vertical axis stands for the relative lost demand compared to that of the oracle (hindsight) scheduling policy. For 4(b), the horizontal axis stands for the value of $h$. Based on different $h$, the scheduling policy outputs different unmet demand, and the vertical axis depicts the ratio of unmet demand to the total demand.